OpenStudy (anonymous):

let f(x)=3x-4 and g(x)=-x^2. Evaluate the composite function. (gf)(0) A.0 B.-16 C.16 D.-10

OpenStudy (jkristia):

what is the resulting function if you have (g o f)(x) ? Replace all 'x' in g(x) with the function defined in f(x) So \[g(x) = x^2 => (f(x))^2 => ??\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

gof=g(f(x)) g(3x-4)=-(3X-4)^2 =-{3(0)-4}^2 putting x=0 =-(-4)^2 =-16

OpenStudy (anonymous):

u r welcome